2. If the test statistic for testing H0: µ≤ 10 vs. HA: µ > 10 at...

2.

If the test statistic for testing H0: µ≤ 10 vs. HA: µ > 10 at the .05 level is 1.641 and the sample size is 20, which of the following is true concerning the correct decision for the test, assuming that σ is unknown?

A Type I error is possible

A Type II error is possible

Either a type I or II error is possible

No error is possible

3. p-value for the test statistic
Answer #1

2) concerning the correct decision type I error is possible in given problem.

3) we are given t test statistic = 1.641 and d.f = n - 1 = 20-1

d.f = 19

p value = P ( t > 1.641 ) at d.f = 19

Using t table

P value = 0.0586